OpenStudy (gabbyalicorn):

What is the value of the ratio of the surface area to the volume of the right prism? A. 1.4 B. 1.07 C. 0.94 D. 0.7

OpenStudy (michele_laino):

here the volume of your prism is: V=6*5*3=...

OpenStudy (gabbyalicorn):

30 x 3 = 60

OpenStudy (michele_laino):

30*3=...

OpenStudy (gabbyalicorn):

90 oops

OpenStudy (michele_laino):

ok! :)

OpenStudy (michele_laino):

whereas the total surface is: A=2*(6*5+6*3+5*3)=...

OpenStudy (gabbyalicorn):

678

OpenStudy (michele_laino):

here is the next step: A=2*(30+18+15)=...

OpenStudy (gabbyalicorn):

126

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

so the requested ratio is: 126/90=...

OpenStudy (gabbyalicorn):

1.4

OpenStudy (michele_laino):

correct! :)

OpenStudy (gabbyalicorn):

:D

OpenStudy (gabbyalicorn):

Thank you!

OpenStudy (michele_laino):

:)